Use Table A.3, Appendix A, to find the following Poisson distribution values.

Appendix AAppendix A Statistical Tables




(Round your answers to 4 decimal places.)

a. P(x = 5 | λ = 1.8) =


b. P(x < 5 | λ = 3.9) =


c. P(x ≥ 3 | λ = 2.5) =


d. P(2 < x ≤ 5 | λ = 4.2) =
Answer #1

Ans:

a)

P(x = 5 | λ = 1.8)=exp(-1.8)*(1.8^5/5!)=0.0260

b)

P(x < 5 | λ = 3.9) = P(x <=4 | λ = 3.9)=Poisson(4,3.9,true)=0.6484

c)

P(x ≥ 3 | λ = 2.5) = 1-P(x <=2 | λ = 2.5)

=1-exp(-2.5)*(2.5^0/0!+2.5^1/1!+2.5^2/2!)

=0.4562

d)

P(2 < x ≤ 5 | λ = 4.2) =exp(-4.2)*(4.2^3/3!+4.2^4/4!+4.2^5/5!)

=0.5429